A certain vibrating string on a piano has a length of 74 cm and forms a standing wave having two antinodes.  

(a) Which harmonic does this wave represents?

(b)  Determine the wavelength of this wave

(c) how many nodes are there if 20.0 Newton find the fundamental frequency I'm the next three frequencies that could cause standing wave patterns on the street



1
Expert's answer
2021-02-19T10:32:03-0500

Answer

a) there is two antinodes then this is third harmonics.

b) it's wavelength is given

"\lambda=\frac{4l}{3}=\frac{4\times0.74}{3}=99.3m"

c) now fundsmental frequency

"f=\frac{v_s}{4l}=\frac{331}{4\times0.74}=111.82Hz"